OpenStudy (anonymous):

Could somebody please walk me through this problem?... (3-3i) divided by / (3+4i)

OpenStudy (anonymous):

Well first you need all like terms together correct? What I would personally do is just make the i values equal

OpenStudy (anonymous):

Wouldn't I do that by multiplying the left side by the right side?

OpenStudy (anonymous):

Like say the equation is this\[\frac{ 3-3i }{3+4i }\]

OpenStudy (anonymous):

It is that (:

OpenStudy (anonymous):

You should try getting the equation to eliminate the i

OpenStudy (anonymous):

So multiply the numerator by the denominator?

OpenStudy (anonymous):

Yes try that

OpenStudy (anonymous):

Multiply the entire division by the denominator so as to eliminate it

OpenStudy (anonymous):

So.... 9-12i?

OpenStudy (anonymous):

Did you foil?

OpenStudy (anonymous):

Ohhh no one second!

OpenStudy (anonymous):

Then it comes out 0?

OpenStudy (anonymous):

Wouldn't the answer be 9-12i?

OpenStudy (anonymous):

No?

OpenStudy (anonymous):

I have options... \[\frac{ (21-21i) }{ 7 }\]

OpenStudy (anonymous):

isnt the equation \[9+12i+9i-12i\]

OpenStudy (anonymous):

\[\frac{ (3-21i) }{ 7 }\]

OpenStudy (anonymous):

\[\frac{ (-3-21i) }{ 25 }\]

OpenStudy (anonymous):

or...

OpenStudy (anonymous):

\[\frac{ (3-21i) }{ 25 }\]

OpenStudy (anonymous):

How are you getting this?

OpenStudy (anonymous):

Listen, after multiplying the bottom of this division by itself it gets eliminated so the bottom is no longer there.

OpenStudy (anonymous):

Then you just multiply the top by what you multiplied the bottom with and there you FOIL

OpenStudy (anonymous):

Those are the options I have to choose from.

OpenStudy (anonymous):

Really? those are the answers?